A seventeen-year, semiannual coupon bond is selling for $1,030. The bond has a face value of $1,000 and a yield to maturity of 5.95 percent. What is the coupon rate?

Multiple Choice

2.65 percent

5.95 percent

6.40 percent

6.23 percent

3.12 percent
Answer #1

M = $1000, n = 17 * 2 = 34 semi-annual periods, i = 5.95%/2 = 2.975% (semi-annual), P = $1030

1030 = C * 21.2074 + 369.0785

660.9215 = C * 21.2074

C = 31.16

Annual Coupon = 2 * $31.16 = $62.33

Coupon Rate = $62.33/$1000 * 100 = 6.23%